# Statistical learning theory and Occam’s razor: The argument from empirical risk minimization

### v1

@@ 5. Conclusion
In this paper, I developed a means-ends model-relative justification for Occam’s razor, based on the method of empirical risk minimization (ERM) in statistical learning theory. I think this account accomodates both the main intuitions and arguments in the machine learning literature in favor of the possibility of such a justification (by actually providing one) and those against (in its various qualifications). It is an honest epistemic justification, that connects a simplicity preference to guarantees of predictive accuracy; and it does not rely on any extra-theoretical assumptions about the true or best hypotheses being simple. But it does presuppose acceptance of the justificatory force of the statistical learning theory framework, including the model-relative nature of the theoretical guarantees; and this model-relativity, I have argued, aligns with the extra-theoretical adoption of a general “forwards-looking” epistemological perspective.
In its essence, the argument capitalizes on a push from the mathematical theory towards specifying a maximally simple hypothesis class. The theoretical push towards simplicity, however, is constrained or indeed opposed by the informal assumptions one brings to the learning problem, which rather pull in the direction of complexity. This means that in those situations in which one’s assumptions do not plausibly allow for a hypothesis class that is sufficiently simple (that still has meaningful theoretical guarantees), the theoretical push towards simplicity, and so the argument, appears moot. However, the reasoning and theoretical results underlying ERM form the core of a more general approach, accommodating weaker assumptions. This approach is automated in the method of structural risk minimization (SRM). It is actually this method that is often brought up in discussions of Occam’s razor, also as representative for a wider family of regularization techniques in machine learning; further, it is closely related to the statistical methods for model selection discussed in the modern philosophical debate. Indeed, to complete the case that the current account accomodates earlier arguments and intuitions pro and con, it needs to be spelled out how exactly it extends to the method of SRM. This I plan to do in a next paper; but I think the main ingredients are already here.

### v2

@@ 5. Conclusion
In this paper, I described a means-ends justificatory argument for Occam’s razor from statistical learning theory, based on the method of empirical risk minimization (ERM). I think this argument accomodates both intuitions and arguments in the machine learning literature in favor of the possibility of such a justification (by actually providing one) and those against (in its various qualifications). It is an honest epistemic justification, that connects a simplicity preference to guarantees of predictive accuracy; and it does not rely on any extra-theoretical assumptions about the true or best hypotheses being simple. But it does presuppose acceptance of the justificatory force of the statistical learning theory framework, including the model-relative nature of the theoretical guarantees. And both the notion of simplicity (as pertaining to classes rather than individual hypotheses) and the means-ends nature of the argument (pushing for simplicity as a constraint on inductive assumptions, rather than directly on inductive conclusions) is perhaps different and more minimal than what one might have hoped from an argument for Occam’s razor.
In its essence, the argument capitalizes on a push from the mathematical theory towards specifying a simple hypothesis class. The theoretical push towards simplicity, however, is checked or indeed opposed by the informal knowledge one brings to the learning problem, which will generally rather pull in the direction of complexity. This points at a certain trade-off, which is in fact automated in the method of structural risk minimization (SRM); and this methods thus appears to directly implement a simplicity preference in inductive inference. It is indeed this method that is usually brought up in discussions of Occam’s razor in statistical learning theory, also as representative for a wider family of regularization techniques in machine learning; further, it is closely related to the statistical methods for model selection discussed in the modern philosophical debate. To complete the case that the current account accomodates earlier arguments and intuitions pro and con, it therefore needs to be spelled out how exactly it figures in the method of SRM. This I intend to do in future work; but I think the core argument is already here.